# Quantum Sensing Using Cold Atoms and Molecules for Nuclear and Particle Physics

### v2

@@ III.1.1 Components of clocks
For a stable reference, it is desirable to have a high sensitivity to small frequency shifts. This requirement can be achieved by using a transition with a large quality factor . Such transitions are forbidden transitions, where a forbidden transition is defined as a transition that cannot be driven by the electric dipole (E1) interaction. The absence of transition probability due to the strongest interaction between atoms and light results in a long-lived excited state, if the excited state is without alternative decay path to the ground state allowed by E1 transitions. In addition, for atomic clocks, it is crucial to select an atomic resonance with minimal sensitivity to external fields.
Historically, microwave transitions between states of two hyperfine levels in the ground state of alkali atoms, such as Cs and Rb, were used as atomic transitions. Over the decades, the accuracy of the Cs atomic clock has improved by six orders of magnitude, as shown in Fig. 2. With the advent of laser spectroscopy, optical atomic clocks have advanced much more rapidly than microwave atomic clocks. One advantage of optical atomic clocks is the transition frequency , which is five orders of magnitude larger than that for microwave atomic clocks and thus helps increase .
Optical clocks are sorted into two types: optical lattice clocks and ion clocks. An optical lattice clock interrogates thousands of atoms trapped in an optical lattice, whereas an ion clock conventionally operates with a single ion trapped in a Paul trap. Initially, the ion clocks were the leading optical clocks in terms of performance. The optical lattice clock was proposed in 2001 [92, 93], and first demonstrated in 2005 [67]. Around this time, the accuracy of optical atomic clocks started to surpass that of microwave clocks, and since then, their accuracies improved by three orders of magnitude. Currently, the smallest fractional uncertainties achieved by optical lattice clocks and ion clocks are comparable, with each type offering distinct advantages and disadvantages. The early development of optical clocks has been summarized well in previous review papers [94, 95].
@@ III.1.2 Allan deviation and the stability of atomic clocks
where is the length of a single measurement cycle, is the total measurement time, is the Wineland parameter for spin squeezing [98], and is the number of atoms in the system [99]. For a coherent spin state [100], , while when the spin state is squeezed in the phase direction during the Ramsey sequence. The optimal is the Ramsay coherence time , which for alkaline-earth-atom-like atoms (see Section III.2.2) can be calculated as [101]
where () is the atom loss rate from the trap for the excited (ground) state, is the Raman scattering rate of the trap laser for the excited state, is the rate at which black body radiation (BBR) induces the transition from the excited state to the state via the state, and is the branching ratio from the to the states.
@@ III.2.1 Alkali-atom-like species
Alkali-atom-like species have a single electron outside a closed shell and have an energy structure shown in Fig. 3(a). A key advantage of these species is the strong transition from the ground state to a excited state suitable for laser cooling. The alkali atoms themselves do not have a narrow-linewidth transition suitable for clock transition, where a clock transition is defined as a transition selected for clock operation. However, some singly-ionized ions of alkaline-earth-atom-like atoms have such narrow-linewidth transitions. In these ions, narrow-linewidth transition used as the clock transition is typically from the ground state to a excited state, corresponding to an electric quadrupole (E2) transition. Examples of ion clocks using this transition are Ca+ [102], Sr+ [103], Ba+ [104], Ra+ [105], and Hg+ [61]. The linewidth of the E2 transition is typically Hz, as listed in Table 1. An exception is Yb+, which also has an electric octupole (E3) transition to an state, in addition to the two E2 transitions to states. Because the E3 transition is a higher-order effect than an E2 transition, it has an exceptionally narrow linewidth of 3.20(16) nHz [106].
@@ III.2.2 Alkaline-earth-atom-like species
Alkaline-earth-atom-like atoms have two valence electrons outside of a closed shell. In the ground state, all electronic spins pair up, resulting in zero electronic angular momentum. Compared to alkali atoms, these species offer a wider range of electronic states, as shown in Fig. 3(b), some of which are advantageous for atomic clocks. The ground state is a singlet state that can be excited to the singlet state via a broad-linewidth E1 transition. This transition is used for the first stage of laser cooling. Among the triplet states, the state has a total angular momentum of , allowing transitions from the ground state that satisfy the selection rule , except that the transition is between a singlet and a triplet state. Such intercombination transitions are weakly allowed due to spin-orbit (L-S) coupling, which results in a significantly narrower linewidth than the transition to the state. The transitions in some atoms, such as Sr and Yb have narrow-linewidth convenient for cooling atoms down to K or lower, leading to the use in the second stage of leaser cooling.
The transition from the ground state to the state is strictly forbidden because both states have . However, if the nucleus has a nonzero magnetic moment, it induces slight mixing between the state and the state, making the transition probability finite. The resulting linewidth is as narrow as mHz, as summarized in Table 1. Also, thanks to the absence of electronic angular momentum, the transition is insensitive to magnetic fields, which is advantageous for stability under ambient fields. Because of these features, the transition is widely used as the clock transition. Note that without a nuclear magnetic moment, a strong magnetic field must be applied to induce mixing at a level comparable to the case with a nuclear magnetic moment. Experimentally, Rabi frequencies of Hz are reported for Sr[107] and Yb[108] with a magnetic field of a few mT and a clock laser intensity of mW/cm2.
@@
The state also has a sufficiently narrow linewidth to serve as a clock transition, because the transition from the ground state to the state is the magnetic quadrupole transition. However, it has not been used as intensively as the state for clock operations. Instead, this state is used as a probe for many-body physics in quantum degenerate gases [116]. Yb has one more metastable state: the state. This state is theoretically predicted to be beneficial for various fundamental physics searches [117, 118]. Recently the direct transitions to this state from the ground state [119, 120] and the state [121] have been observed for the first time. Although the uncertainty in its absolute frequency measurement still needs improvement before it can be used as a clock transition, an initial analysis of its potential application in nuclear and particle physics has already been reported [122].
@@ III.3 Optical lattice clock
The stability of the state-of-the-art optical lattice clock reached in 2013 [132]. In this report, the stability at 1 s was , and by integrating over s, the stability reached the low range. This high stability is achieved by a combination of low QPN of with and ms. to further improve the SQL, increasing is crucial, which requires improving the stability of the clock laser. Efforts to develop more stable clock lasers are ongoing using cryogenic cavities [133, 134], achieving a fractional stability of from 0.8 s to 80 s, or 16 mHz linewidth for 1542 nm laser.
A recently developed technique for quickly estimating systematic shifts is the differential measurement of multiple atomic ensembles with a common clock laser. The advantage of this method is that the differential response of atoms is immune to laser fluctuation and noise, including the Dick effect [135, 136], enabling precise comparisons even with a less stable laser. This method can be implemented various ways, such as two atomic ensembles in separate vacuum chambers[137], multiple atomic ensembles in the same vacuum chamber[138], and a single atomic ensemble analyzed in two subregions using camera imaging [139]. The highest reported stability at 1 s is , while the record long-term stability of an atomic ensemble is [139]. Higher stability allows for the estimation of systematic uncertainties in a shorter time, contributing to improving clock accuracy. An optical lattice clock with a fractional uncertainty of was reported in 2015 [75], and now the best performing clock has a fractional uncertainty of [78].
@@ III.3.1 Major systematic uncertainties
For an optical clock to be highly accurate, it is essential to accurately estimate systematic shifts. Optical lattice clocks are typically stable enough that statistical uncertainty can be minimized by increasing the overall measurement time. As a result, their overall uncertainties are primarily limited by systematic uncertainties. Historically, improvements in the accuracy of optical lattice clocks have been enabled by a deeper understanding of systematic shifts. The pursuit of better uncertainty estimates have motivated the development of techniques to further mitigate these shifts. A notable example is the dc Stark shift. When optical lattice clocks were reported for the first time, the dc Stark shift caused by stray electric fields in a standard vacuum chamber were not a major contributor to total systematic uncertainty. However, as the overall uncertainty of optical lattice clocks improved, the impact of ambient electric fields became increasingly significant[140]. To precisely quantify this effect, the dc Stark shift was carefully measured with a chamber with electrodes inside [141]. In addition, to prevent charge accumulation on insulator surfaces, the inner surface of the viewports has conductive coating, typically indium tin oxide, which helped suppress stray electric fields.
One of the major sources of uncertainty in optical lattice clock is the BBR shift. This uncertainty arises from two factors: the amount of BBR experienced by the atoms and the uncertainty in the polarizability of the ground and particularly the excited state. To estimate the polarizability precisely, the transition matrix for the transition, which has the smallest transition frequency among the E1 transitions from the state, was carefully measured for Sr [78] and Yb [142]. To control the amount of BBR, a metallic shield with stabilized temperature is implemented inside the vacuum chamber. The temperature is calibrated to the 1-mK level for Sr[75] and Yb[125], confirmed by a Pt thermometer located near the atoms. In addition, the inner surfaces of both the vacuum chamber and the shield are coated with carbon nanotube to keep the emissivity of the chamber wall as close to 1 as possible [143]. Surrounding atoms with a cold shield inside a vacuum chamber reduces the amount of BBR at the atom location, as dictated by the Stefan-Boltzmann law. Such a cryogenic optical lattice clock at 95 K achieved a small BBR uncertainty [74].
@@
Another major source of uncertainty originates from the ac Stark shift caused by the trapping laser (Ref.[78] for Sr, [125] for Yb). When the total uncertainty was on the order of , the shift was mitigated by carefully selecting magic wavelength for Sr[70] and Yb[72]. However, with an orders of magnitude smaller uncertainty, higher order effects must also be taken into account. One such effect arises from vector and tensor shifts, which depends on the polarizations of the trapping lasers[146]. The shift proportional to the trap depth is formulated as
where , with , , and being the lattice polarization vector, lattice propagation vector, and lattice polarization ellipticity and being the direction of the quantization axis. , , and are the differential scalar, vector, and tensor shift coefficients. For a one-dimensional lattice, properly selecting the relative orientation between the polarization and the magnetic field can cancel some of these shifts. However, in a three-dimensional lattice, such cancellation is not possible in all three dimensions. This limitation is one of the reasons why state-of-the-art optical lattice clocks use one-dimensional optical lattice.
@@
where with being recoil energy for lattice light, is the quantum number for the longitudinal vibrational state, is the differential E1 polarizability between the ground and excited states of the clock transition, is the lattice laser frequency, is the detuning of the lattice laser frequency from the E1 magic frequency and and are the differential multipolarizability and hyperpolarizability, respectively. This formulation accounts for the effect of the longitudinal temperature of atoms in the lattice via and incorporates the shifts in Eq. 4. The radial temperature can be included by modifying as , where is the exponent on for each term in Eq. III.3.1. Using this formalism, the concept of operational magic intensity was itroduced [147]. However, despite these characterizations, the lattice ac Stark shift remains one of the major sources of uncertainty in the best optical lattice clocks. A careful analysis on the order of is reported for Sr [148], leading to a fractional uncertainty of [78]. Similar analysis is also going on for Yb [149]. An alternative approach to suppress the uncertainty from the ac Stark shift is to use atoms with small higher-order effects. Particularly, hyperpolarizability for Hg, Cd, and Zn is calculated to be smaller than for Sr and Yb [150]. These clocks can be potentially useful to suppress uncertainties related to the lattice ac Stark shift, in addition to the BBR shift.
@@
The density shift, which arises from atom-atom interactions, is another source of uncertainty. Since the shift is proportional to the density of atoms in the lattice, reducing the density helps mitigate its effects. Another way to control the density shift is by trapping atoms in a three-dimensional optical lattice and removing excess atoms by lowering the Fermi temperature of a degenerate Fermi gas[151]. However, both approaches have drawbacks. Reducing the density of atoms increases the QPN due to the small number of atoms, while the three-dimensional optical lattice induces a large uncertainty in the ac Stark shift due to the tensor shift. Additionally, the edge effect [152] needs to be carefully managed. Recent studies utilize low-density fermionic gas in large vertical one-dimensional optical lattices. First, a large optical lattice, e.g., a beam waist size of 260 m [139], is used to reduce the density of atoms. To achieve such a lattice, trap laser intensity is enhanced with an intermediate finesse cavity ([153], [139]). Second, fermionic isotopes naturally suppress the collisional shift by s-wave collisions. For Sr, the remaining shift due to p-wave collisions is canceled by inter-layer s-wave collisions, which is induced by the delocalized atoms over a few layers due to the low trap depth () [139]. This approach has enabled the clock to reach fractional uncertainty [78]. For Yb, delocalization of atomic cloud to reduce the density in this large one-dimensional optical lattice is performed. [154] Other systematic uncertainties, such as background gas collision shift, second order Zeeman shift, and servo error are well-controlled and significantly suppressed compared to the major uncertainties discussed here.

### v3

@@ III.1.1 Components of clocks
For a stable reference, it is desirable to have a high sensitivity to small frequency shifts. This requirement can be achieved by using a transition with a large quality factor . Such transitions are forbidden transitions, where a forbidden transition is defined as a transition that cannot be driven by the electric dipole (E1) interaction. The absence of transition probability due to the strongest interaction between atoms and light results in a long-lived excited state, if the excited state is without an alternative decay path to the ground state allowed by E1 transitions. In addition, for atomic clocks, it is crucial to select an atomic resonance with minimal sensitivity to external fields.
Historically, microwave transitions between states of two hyperfine levels in the ground state of alkali atoms, such as Cs and Rb, were used as the atomic transitions. Over the decades, the accuracy of the Cs atomic clock has improved by six orders of magnitude, as shown in Fig. 2. With the advent of laser spectroscopy, optical atomic clocks have advanced much more rapidly than microwave atomic clocks. One advantage of optical atomic clocks is the transition frequency , which is five orders of magnitude larger than that for microwave atomic clocks and thus helps increase .
Optical atomic clocks are sorted into two types: optical lattice clocks and ion clocks. An optical lattice clock interrogates thousands of atoms trapped in an optical lattice, whereas an ion clock conventionally operates with a single ion trapped in a Paul trap. Initially, ion clocks were the leading optical atomic clocks in terms of performance. The optical lattice clock was proposed in 2001 [95, 96], and first demonstrated in 2005 [70]. Around this time, the accuracy of optical atomic clocks started to surpass that of microwave atomic clocks, and since then, their accuracies improved by three orders of magnitude. Currently, the smallest fractional uncertainties achieved by optical lattice clocks and ion clocks are comparable, with each type offering distinct advantages and disadvantages. The early development of optical atomic clocks has been summarized well in previous review papers [97, 98].
@@ III.1.2 Allan deviation and the stability of atomic clocks
where is the length of a single measurement cycle, is the total measurement time, is the Wineland parameter for spin squeezing [101], and is the number of atoms in the system [102]. For a coherent spin state [103], , while when the spin state is squeezed in the phase direction during the Ramsey sequence.
@@ III.2.1 Alkali-atom-like species
Alkali-atom-like species have a single electron outside a closed shell and have an energy structure shown in Fig. 3(a). A key advantage of these species is the strong transition from the ground state to a excited state suitable for laser cooling. The alkali atoms themselves do not have a narrow-linewidth transition suitable for clock transition, where a clock transition is defined as a transition selected for clock operation. However, some singly-ionized ions of alkaline-earth-atom-like atoms have such narrow-linewidth transitions. In these ions, narrow-linewidth transition used as the clock transition is typically from the ground state to a excited state, corresponding to an electric quadrupole (E2) transition. Examples of ion clocks using this transition are Ca+ [104], Sr+ [105], Ba+ [106], Ra+ [107], and Hg+ [61]. The linewidth of the E2 transition is typically Hz, as listed in Table 1. An exception is Yb+, which also has an electric octupole (E3) transition to an state, in addition to the two E2 transitions to states. Because the E3 transition is a higher-order effect than an E2 transition, it has an exceptionally narrow linewidth of nHz [108].
@@ III.2.2 Alkaline-earth-atom-like species
Alkaline-earth-atom-like atoms have two valence electrons outside of a closed shell. In the ground state, all electronic spins pair up, resulting in zero electronic angular momentum. Compared to alkali atoms, these species offer a wider range of electronic states, as shown in Fig. 3(b), some of which are advantageous for atomic clocks. The ground state is a singlet state that can be excited to the singlet state via a broad-linewidth E1 transition. This transition is used for the first stage of laser cooling. Among the triplet states, the state has a total angular momentum of , allowing transitions from the ground state that satisfy the selection rule , except that the transition is between a singlet and a triplet state. Such intercombination transitions are weakly allowed due to spin-orbit (L-S) coupling, which results in a significantly narrower linewidth than the transition to the state. The transitions in some atoms, such as Sr and Yb have narrow linewidth convenient for cooling atoms down to K or lower, leading to the use in the second stage of laser cooling.
The transition from the ground state to the state is strictly forbidden because both states have . However, if the nucleus has a nonzero magnetic moment, it induces slight mixing between the state and the state, making the transition probability finite. The resulting linewidth is as narrow as mHz, as summarized in Table 1. Also, thanks to the absence of electronic angular momentum, the transition is insensitive to magnetic fields, which is advantageous for stability under ambient fields. Because of these features, the transition is widely used as the clock transition. Note that without a nuclear magnetic moment, a strong magnetic field must be applied to induce mixing at a level comparable to the case with a nuclear magnetic moment [109]. Experimentally, Rabi frequencies of Hz are reported for Sr[110] and Yb[111] with a magnetic field of a few mT and a clock laser intensity of mW/cm2.
@@
The state also has a sufficiently narrow linewidth to serve as a clock transition, because the transition from the ground state to the state is the magnetic quadrupole transition. However, it has not been used as intensively as the state for clock operations. Instead, this state is used as a probe for many-body physics in quantum degenerate gases [119]. Yb has one more metastable state: the state. This state is theoretically predicted to be beneficial for various fundamental physics searches [120, 121]. Recently the direct transitions to this state from the ground state [122, 123] and the state [124] have been observed for the first time. Although the uncertainty in its absolute frequency measurement still needs improvement before it can be used as a clock transition, an initial analysis of its potential application in nuclear and particle physics has already been published [125].
@@ III.3 Optical lattice clock
The stability of the state-of-the-art optical lattice clock reached in 2013 [134]. In this report, the stability at 1 s was , and by integrating over s, the stability reached the low range. This high stability is achieved by a combination of low QPN of with and ms. To further improve the SQL, increasing is crucial, which requires improving the stability of the clock laser. Efforts to develop more stable clock lasers are ongoing using cryogenic cavities [135, 136], achieving a fractional stability of from 0.8 s to 80 s, or 16 mHz linewidth for 1542 nm laser. Another cryogenic silicon cavity for 1397 nm laser reached a stability of from 4 to 12 s, corresponding to a linewidth of 9.6(3) mHz linewidth [137].
A recently developed technique for quickly estimating systematic shifts is the differential measurement of multiple atomic ensembles with a common clock laser. The advantage of this method is that the differential response of atoms is immune to laser fluctuation and noise, including the Dick effect [138, 139], enabling precise comparisons even with a less stable laser. This method can be implemented various ways, such as two atomic ensembles in separate vacuum chambers[140], multiple atomic ensembles in the same vacuum chamber[141], and a single atomic ensemble analyzed in two subregions using camera imaging [142]. The highest reported stability at 1 s is [143], while the record long-term stability of an atomic ensemble is [142]. Higher stability allows for the estimation of systematic uncertainties in a shorter time, contributing to improving clock accuracy. An optical lattice clock with a fractional uncertainty of was reported in 2015 [78], and now the best performing clock has a fractional uncertainty of [81].
This method of comparing multiple atomic ensembles is used to investigate higher atomic coherence time. Best coherence time achieved is longer than 100 s. Ref. [143] made use of low-density atoms trapped in a shallow optical lattice to suppress the decoherence due to the Raman scattering of the excited state by the lattice light and atomic collision. Ref. [144] utilized hyperfine-resolved readout to enhance the coherence.
@@ III.3.1 Major systematic uncertainties
For an optical atomic clock to be highly accurate, it is essential to accurately estimate systematic shifts. Optical lattice clocks are typically stable enough that statistical uncertainty can be minimized by increasing the overall measurement time. As a result, their overall uncertainties are primarily limited by systematic uncertainties. Historically, improvements in the accuracy of optical lattice clocks have been enabled by a deeper understanding of systematic shifts. The pursuit of better uncertainty estimates have motivated the development of techniques to further mitigate these shifts. A notable example is the dc Stark shift. When optical lattice clocks were reported for the first time, the dc Stark shift caused by stray electric fields in a standard vacuum chamber were not a major contributor to total systematic uncertainty. However, as the overall uncertainty of optical lattice clocks improved, the impact of ambient electric fields became increasingly significant[145]. To precisely quantify this effect, the dc Stark shift was carefully measured with a chamber with electrodes inside [146]. In addition, to prevent charge accumulation on insulator surfaces, the inner surface of the viewports has conductive coating, typically indium tin oxide, which helped suppress stray electric fields.
One of the major sources of uncertainty in optical lattice clock is the black body radiation (BBR) shift. This uncertainty arises from two factors: the amount of BBR experienced by the atoms and the uncertainty in the polarizability of the ground and, particularly, the excited state. To estimate the polarizability precisely, the transition matrix for the transition, which has the smallest transition frequency among the E1 transitions from the state, was carefully measured for Sr [81] and Yb [147]. To control the amount of BBR, a metallic shield with stabilized temperature is implemented inside the vacuum chamber. The temperature is calibrated to the 1-mK level for Sr[78] and Yb[79], confirmed by a Pt thermometer located near the atoms. In addition, the inner surfaces of both the vacuum chamber and the shield are coated with carbon nanotube to keep the emissivity of the chamber wall as close to 1 as possible [148]. Surrounding atoms with a cold shield inside a vacuum chamber reduces the amount of BBR at the atom location, as dictated by the Stefan-Boltzmann law. Such a cryogenic optical lattice clock for Sr at 95 K achieved a small BBR uncertainty [77], with thorough uncertainty budget listed. A Yb optical lattice clock with a cryogenic shield is recently reported [149]. The temperature is set at 77 K, and achieved fractional uncertainty for the BBR shift is .
@@
Another major source of uncertainty originates from the ac Stark shift caused by the trapping laser (Ref.[81] for Sr, Ref.[79] for Yb). When the total uncertainty was on the order of , the shift was mitigated by carefully selecting magic wavelength for Sr[73] and Yb[75]. However, with an orders of magnitude smaller uncertainty, higher order effects must also be taken into account. One such effect arises from vector and tensor shifts, which depends on the polarizations of the trapping lasers[152]. The shift proportional to the trap depth is formulated as
where , with , , and being the lattice polarization vector, lattice propagation vector, and lattice polarization ellipticity, respectively, and being the direction of the quantization axis. , , and are the differential scalar, vector, and tensor shift coefficients, respectively. For a one-dimensional lattice, properly selecting the relative orientation between the polarization and the magnetic field can cancel some of these shifts. However, in a three-dimensional lattice, such cancellation is not possible in all three dimensions. This limitation is one of the reasons why state-of-the-art optical lattice clocks use a one-dimensional optical lattice.
@@
where with being recoil energy for lattice light, is the differential E1 polarizability between the ground and excited states of the clock transition, is the lattice laser frequency, is the detuning of the lattice laser frequency from the E1 magic frequency and and are the differential multipolarizability and hyperpolarizability, respectively. This formulation accounts for the effect of the longitudinal temperature of atoms in the lattice via and incorporates the shifts in Eq. 3. The radial temperature can be included by modifying as , where is the exponent on for each term in Eq. III.3.1. Using this formalism, the concept of operational magic intensity was itroduced [153]. However, despite these characterizations, the lattice ac Stark shift remains one of the major sources of uncertainty in the best optical lattice clocks. A careful analysis on the order of is reported for Sr [154], leading to a fractional uncertainty of [81]. Similar analysis is also going on for Yb [155]. An alternative approach to suppress the uncertainty from the ac Stark shift is to use atoms with small higher-order effects. Particularly, hyperpolarizability for Hg, Cd, and Zn is calculated to be smaller than that for Sr and Yb [156]. These clocks can be potentially useful to suppress uncertainties related to the lattice ac Stark shift, in addition to the BBR shift.
@@
The density shift, which arises from atom-atom interactions, is another source of uncertainty. Since the shift is proportional to the density of atoms in the lattice, reducing the density helps mitigate its effects. Another way to control the density shift is by trapping atoms in a three-dimensional optical lattice and removing excess atoms by lowering the Fermi temperature of a degenerate Fermi gas[157]. However, both approaches have drawbacks. Reducing the density of atoms increases the QPN due to the small number of atoms, while the three-dimensional optical lattice induces a large uncertainty in the ac Stark shift due to the tensor shift. Additionally, the edge effect [158] needs to be carefully managed. Recent studies utilize low-density fermionic gas in large vertical one-dimensional optical lattices. First, a large optical lattice, e.g., a beam waist size of 260 m [142], is used to reduce the density of atoms. To achieve such a lattice, trap laser intensity is enhanced with an intermediate finesse cavity ([159], [142]). Second, fermionic isotopes naturally suppress the collisional shift by s-wave collisions. For Sr, the remaining shift due to p-wave collisions is canceled by inter-layer s-wave collisions, which are induced by the delocalized atoms over a few layers due to the low trap depth () [142]. This approach has enabled the clock to reach fractional uncertainty [81]. For Yb, delocalization of the atomic cloud to reduce the density in this large one-dimensional optical lattice is performed. [160] Other systematic uncertainties, such as the background gas collision shift, the second order Zeeman shift, and the servo error are well-controlled and significantly suppressed compared to the major uncertainties discussed here.